# Strange stars admixed with dark matter: equiparticle model in a two fluid approach

### v1

@@ I Introduction
Recently many researchers have been dedicated efforts to better comprehend the physics of these objects, that may present another mysterious component as well, named as dark matter (DM) [17, 18]. The main evidences for the existence of nonluminous matter come from the analysis of galaxies rotation curves, and from the use of gravitational lensing methods [19, 20], that lead to visible galaxy clusters masses of around 10% to 20%, being the remaining mass due to DM. Measurements of the anisotropy of cosmic microwave background also suggest presence of DM [21, 22, 23, 24, 25, 26]. It is a consensus that the universe, as we know nowadays, is 27% composed by DM, 68% of dark energy, and 5% of ordinary matter.
@@ III Dark matter models
with $C_{V}=g_{V}/m_{V}$ and $\rho_{\chi}=k_{F\chi}^{3}/(3\pi^{2})$. The Fermi momentum of the dark particle is $k_{F\chi}$.
@@
with $C_{\sigma\phi}=g_{\sigma}/m_{\phi}$. The DM density $\rho_{\sigma}$ relates to the zero component of the vector field $\phi_{0}$, and with the scalar field and its conjugate through $\phi_{0}=(g_{\sigma}/m_{\phi}^{2})\rho_{\sigma}$, and $\rho_{\sigma}=2m_{\sigma}\sigma^{*}\sigma$, respectively. Detailed derivation of these expressions can be found in [77].
@@ IV Strange stars admixed with dark matter
Here we impose that every star present a certain fraction of dark matter in its composition, i.e., a fixed value of
$F_{\mbox{\tiny DM}}=\frac{M_{\mbox{\tiny DM}}}{M}.$ (25)
In order to satisfy this restriction, we proceed as follows. First one needs to construct all the inputs for the TOV equations, namely, $P^{c}_{\mbox{\tiny vis}}$, $\mathcal{E}^{c}_{\mbox{\tiny vis}}$, $P^{c}_{\mbox{\tiny DM}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$ (despite not directly used as initial conditions, the energy densities have also to be furnished since their relationships with the pressures are used to replace $\mathcal{E}$’s by $P$’s in the TOV equations). The visible matter equation of state (EOS) is straightforward, directly evaluated from Eqs. (11)-(12). For both DM models, fermionic and bosonic, we connect them with the EQP model by imposing that $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ where the energy density fraction, $f$, is a constant [70]. In the case of the bosonic DM model, this leads to an analytical expression for the dark boson density,
@@
Since the connection $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ is established, we solve the two-fluid TOV equations for huge set of values for $f$. For each solution related to a particular $f$, we evaluate Eq. (25) and select, among all the solutions, only the $(M,R)$ pairs that satisfy the particular value of $F_{\mbox{\tiny DM}}$ chosen. The final result consists in a mass-radius diagram representing quark stars, each one of them admixed with a fixed fraction of dark matter.
@@
In the case of the two-fluid method, the generalization determines that at the onset of unstable configurations the number of visible ($N_{\mbox{\tiny vis}}$) and dark ($N_{\mbox{\tiny DM}}$) particles are stationary under variations of $\mathcal{E}^{c}_{\mbox{\tiny vis}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$. As pointed out in [81, 80], it is analogous to diagonalize the matrix $\partial N_{i}/\partial\mathcal{E}^{c}_{j}$ ($i,j$=VIS, DM) with the aim of searching for its associated eingenvalues, claimed to be both positive for stable stars configurations. However, as discussed in [82], there is an indication that DM models presenting central pressures in the range of $P^{c}_{\mbox{\tiny DM}}\lesssim 5.1$ fm-4 generate stars in a stable region. Both DM models used in this work, bosonic and fermionic ones, are in agreement with this constraint. Furthermore, the authors of [83, 84] have used in their calculations, based on previous studies [85, 86], that the maxima of the fixed $F_{\mbox{\tiny DM}}$ curves represent the last stable stars in the mass-radius diagram, as it is in the case of one fluid systems. Here we adopt the same procedure and exclude points beyond the maximum one. The final mass-radius diagrams for both models described here are shown in Fig. 3.
@@
It is worth noticing the strong linear relation between the maximum mass and the dark matter content. As a consequence, one has a constant decreasing rate of $M_{\mbox{\tiny max}}$ as a function of $F_{\mbox{\tiny DM}}$. Such behavior is hard to be predicted before calculations due to the nontrivial structure of both, the two-fluid equations and the equations of state used as input, in our case, the EQP model in the visible sector, and the fermionic/bosonic models in the description of dark matter. We also furnish the slope of the linear fitting obtained, namely, $-0.0342$ ($-0.315$) for bosonic (fermionic) model. Notice that both are very similar. Furthermore, these values are very close to $-0.0362$, result found in [92] in a study performed with a completely different quark model in the ordinary matter sector: the MIT bag model. For the dark sector, they have used a model made up of fermionic particles of mass $100$ GeV, number more than $50$ times greater than the value used in the current work for the FDM model. This result maybe indicates a possible universality in the (linear) decreasing behavior of the maximum mass in quark models admixed with dark matter.
@@
Still regarding the features related to Fig. 3, we emphasize that the difference between the entire mass-radius curves generated from the quark-FDM approach and from the quark-BDM one are very small. We assign this effect to the dominance of the effective quark model over both, bosonic and fermionic DM models used in the analysis. In order to illustrate this statement, we display in Figs. 5 the profiles of the same star constructed by using both BDM and FDM models, with $M=1.806M_{\odot}$ and $10\%$ of dark matter in its composition.
@@
Figure 5: Pressure (panels a and b) and enclosed mass (panels c and d) as functions of radial coordinate for a mass fraction equal to $10\%$ and total stellar mass of $M=1.806M_{\odot}$. Full (dashed) lines: visible (dark) matter. For the bosonic (fermionic) case in panel a (b), the energy density fraction is $f=0.87$ ($0.6$).
@@
Notice, from this example, that the radius of the star is $R=R_{\mbox{\tiny vis}}\simeq 13.6$ km, since $R_{\mbox{\tiny vis}}>R_{\mbox{\tiny DM}}$ in this case, independent on the model used for the dark sector. The only modification caused by the interchange of the DM model is the central pressure (at $r=0$) of both sectors. Nevertheless, $R_{\mbox{\tiny vis}}$ is still pratically equal for both models and determines the radius of the final star, according to the criterion used to define this quantity ($R$ corresponds to the outermost radius between $R_{\mbox{\tiny vis}}$ and $R_{\mbox{\tiny DM}}$). However, the situation is not the same for stars in which a DM halo is formed around it. We can verify such an effect for the case of $M=1.704M_{\odot}$ and $F_{\mbox{\tiny DM}}=9\%$, depicted in Fig. 6, for the strange star constructed with inclusion of bosonic dark matter in particular.
@@
Figure 6: Pressure as a function of radial coordinate for a mass fraction equal to $9\%$ and total stellar mass of $M=1.704M_{\odot}$. Full (dashed) lines: visible (dark) matter. Strange star admixed with bosonic dark matter with $m_{\sigma}=15$ GeV (panel a) and $m_{\sigma}=2$ GeV (panel b).
@@
In this figure we present two parametrizations of the BDM model, specifically with different values of the dark boson mass, namely, $m_{\sigma}=15$ GeV (standard value previously used) and $m_{\sigma}=2$ GeV. Notice that the decrease of the dark particle mass induces the situation of $R_{\mbox{\tiny DM}}>R_{\mbox{\tiny vis}}$ producing, consequently, the formation of a DM halo. In this case, the radius of the final star changes to $R=R_{\mbox{\tiny vis}}\simeq 13.9$ km to $R=R_{\mbox{\tiny DM}}\simeq 16.3$ km. Since fermionic and bosonic models generate similar mass-radius diagrams for $m_{\sigma}=15$ GeV and $m_{\chi}=1.9$ GeV, the decrease of $m_{\sigma}$ will produce a more pronounced difference between the diagrams obtained from FDM model with $m_{\chi}=1.9$ GeV in comparison with the ones determined from BDM model with $m_{\sigma}=2$ GeV, due to the emergence of DM halos in the latter system. These features are compatible with studies presented, for instance, in [91] where the authors used the two-fluid approach with bosonic DM, and in [93] where a fermionic DM model is applied in the description of stellar matter. In these works, the DM halo formation is observed when the dark particle (bosons or fermions) are decreased. In particular, since they treat neutron stars, hadronic models are used in the visible matter sector. In our work, on the other hand, strange stars are studied but the same effect is exhibited as well.

### v2

@@ I Introduction
Recently many researchers have been dedicated efforts to better comprehend the physics of these objects, that may present another mysterious component as well, named as dark matter (DM) [17, 18, 19]. The main evidences for the existence of non-luminous matter come from the analysis of galaxies rotation curves, and from the use of gravitational lensing methods [20, 21], that lead to visible galaxy clusters masses of around 10% to 20%, being the remaining mass due to DM. Measurements of the anisotropy of cosmic microwave background also suggest presence of DM [22, 23, 24, 25, 26, 27]. It is a consensus that the universe, as we know nowadays, is 27% composed by DM, 68% of dark energy, and 5% of ordinary matter.
@@ III Dark matter models
with $C_{V}=g_{V}/m_{V}$ and $\rho_{\chi}=k_{F\chi}^{3}/(3\pi^{2})$. The Fermi momentum of the dark particle is $k_{F\chi}$. Fermionic self-interacting dark matter models coupled to hadronic matter were used in the description of neutron stars with DM content, as the reader can find in [74], for instance.
@@
with $C_{\sigma\phi}=g_{\sigma}/m_{\phi}$. The DM density $\rho_{\sigma}$ relates to the zero component of the vector field $\phi_{0}$, and with the scalar field and its conjugate through $\phi_{0}=(g_{\sigma}/m_{\phi}^{2})\rho_{\sigma}$, and $\rho_{\sigma}=2m_{\sigma}\sigma^{*}\sigma$, respectively. Detailed derivation of these expressions can be found in [79]. We also address the reader to [80] for a deeper study of scalar fields, in general, in the gravitational context. Other similar bosonic DM models coupled to luminous matter, studying neutron stars in particular, were analyzed in [81, 82].
@@ IV Strange stars admixed with dark matter
Here we impose that every star present a certain fraction of dark matter in its composition, i.e., a fixed value of $F_{\mbox{\tiny DM}}=M_{\mbox{\tiny DM}}/M$. In order to satisfy this restriction, we proceed as follows. First one needs to construct all the inputs for the TOV equations, namely, $P^{c}_{\mbox{\tiny vis}}$, $\mathcal{E}^{c}_{\mbox{\tiny vis}}$, $P^{c}_{\mbox{\tiny DM}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$ (despite not directly used as initial conditions, the energy densities have also to be furnished since their relationships with the pressures are used to replace $\mathcal{E}$’s by $P$’s in the TOV equations). The visible matter equation of state (EOS) is straightforward, directly evaluated from Eqs. (11)-(12). For both DM models, fermionic and bosonic, we connect them with the EQP model by imposing that $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ where the energy density fraction, $f$, is a constant [71]. In the case of the bosonic DM model, this leads to an analytical expression for the dark boson density,
@@
Since the connection $\mathcal{E}_{\mbox{\tiny DM}}=f\mathcal{E}_{\mbox{\tiny vis}}$ is established, we solve the two-fluid TOV equations for huge set of values for $f$. For each solution related to a particular $f$, we evaluate $F_{\mbox{\tiny DM}}$ and select, among all the solutions, only the $(M,R)$ pairs that satisfy the particular value of $F_{\mbox{\tiny DM}}$ chosen. The final result consists in a mass-radius diagram representing quark stars, each one of them admixed with a fixed fraction of dark matter.
@@
In the case of the two-fluid method, the generalization determines that at the onset of unstable configurations the number of visible ($N_{\mbox{\tiny vis}}$) and dark ($N_{\mbox{\tiny DM}}$) particles are stationary under variations of $\mathcal{E}^{c}_{\mbox{\tiny vis}}$ and $\mathcal{E}^{c}_{\mbox{\tiny DM}}$. As pointed out in [86, 85], it is analogous to diagonalize the matrix $\partial N_{i}/\partial\mathcal{E}^{c}_{j}$ ($i,j$=VIS, DM) with the aim of searching for its associated eingenvalues, claimed to be both positive for stable stars configurations. However, as discussed in [87], there is an indication that DM models presenting central pressures in the range of $P^{c}_{\mbox{\tiny DM}}\lesssim 5.1$ fm-4 generate stars in a stable region. Both DM models used in this work, bosonic and fermionic ones, are in agreement with this constraint. Furthermore, the authors of [88, 89] have used in their calculations, based on previous studies [90, 91], that the maxima of the fixed $F_{\mbox{\tiny DM}}$ curves represent the last stable stars in the mass-radius diagram, as it is in the case of one fluid systems. In our study we adopt the same procedure and exclude points beyond the maximum one. The final mass-radius diagrams for both models described here are shown in Fig. 3.
@@
It is worth noticing the strong linear relation between the maximum mass and the dark matter content. As a consequence, one has a constant decreasing rate of $M_{\mbox{\tiny max}}$ as a function of $F_{\mbox{\tiny DM}}$. Such behavior is hard to be predicted before calculations due to the nontrivial structure of both, the two-fluid equations and the equations of state used as input, in our case, the EQP model in the visible sector, and the fermionic/bosonic models in the description of dark matter. We also furnish the slope of the linear fitting obtained, namely, $-0.0343$ ($-0.315$) for bosonic (fermionic) model. Notice that both are very similar. Furthermore, these values are very close to $-0.0362$, result found in [97] in a study performed with a completely different quark model in the ordinary matter sector: the MIT bag model. For the dark sector, they have used a model made up of fermionic particles of mass $100$ GeV, number more than $50$ times greater than the value used in the current work for the FDM model. This result maybe indicates a possible universality in the (linear) decreasing behavior of the maximum mass in quark models admixed with dark matter.
@@
Still regarding the features related to Fig. 3, we emphasize that the difference between the entire mass-radius curves generated from the quark-FDM approach and from the quark-BDM one are very small. We assign this effect to the dominance of the effective quark model over both, bosonic and fermionic DM models used in the analysis. In order to illustrate this statement, we display in Figs. 5 the profiles of the same star constructed by using both BDM and FDM models, with $M=1.81M_{\odot}$ and $10\%$ of dark matter in its composition.
@@
Figure 5: Pressure (panels a and b) and enclosed mass (panels c and d) as functions of radial coordinate for a mass fraction equal to $10\%$ and total stellar mass of $M=1.81M_{\odot}$. Full (dashed) lines: visible (dark) matter. For the bosonic (fermionic) case in panel a (b), the energy density fraction is $f=0.87$ ($0.6$).
@@
Notice from this example that the radius of the star is $R=R_{\mbox{\tiny vis}}\simeq 13.6$ ($13.7$) km, since $R_{\mbox{\tiny vis}}>R_{\mbox{\tiny DM}}$ in this case, by using the bosonic (fermionic) model in the dark sector. The main modification caused by the interchange of the DM model is the central pressure (at $r=0$) in both sectors. Nevertheless, $R_{\mbox{\tiny vis}}$ is still pratically equal for both models and determines the radius of the final star, according to the criterion used to define this quantity ($R$ corresponds to the outermost radius between $R_{\mbox{\tiny vis}}$ and $R_{\mbox{\tiny DM}}$). However, the situation is not the same for stars in which a DM halo is formed around it. We can verify such an effect for the case of $M=1.86M_{\odot}$ and $F_{\mbox{\tiny DM}}=10\%$, depicted in Fig. 6, for the strange star constructed with inclusion of bosonic dark matter in particular.
@@
Figure 6: Pressure as a function of radial coordinate for a mass fraction equal to $10\%$ and total stellar mass of $M=1.86M_{\odot}$. Full (dashed) lines: visible (dark) matter. Strange star admixed with bosonic dark matter with $m_{\sigma}=15$ GeV (panel a) and $m_{\sigma}=2$ GeV (panel b).
@@
In this figure we present two parametrizations of the BDM model, specifically with different values of the dark boson mass, namely, $m_{\sigma}=15$ GeV (standard value previously used) and $m_{\sigma}=2$ GeV. Notice that the decrease of the dark particle mass induces the situation of $R_{\mbox{\tiny DM}}>R_{\mbox{\tiny vis}}$ producing, consequently, the formation of a DM halo. In this case, the radius of the final star changes from $R=R_{\mbox{\tiny vis}}\simeq 13$ km to $R=R_{\mbox{\tiny DM}}\simeq 19$ km. Since fermionic and bosonic models generate similar mass-radius diagrams for $m_{\sigma}=15$ GeV and $m_{\chi}=1.9$ GeV, the decrease of $m_{\sigma}$ will produce a more pronounced difference between the diagrams obtained from FDM model with $m_{\chi}=1.9$ GeV in comparison with the ones determined from BDM model with $m_{\sigma}=2$ GeV, due to the emergence of DM halos in the latter system. These features are compatible with studies presented, for instance, in [96] where the authors used the two-fluid approach with bosonic DM, and in [98] where a fermionic DM model is applied in the description of stellar matter. In these works, the DM halo formation is observed when the dark particle (bosons or fermions) are decreased. In particular, since they treat neutron stars, hadronic models are used in the visible matter sector. In our work, on the other hand, strange stars are studied but the same effect is exhibited as well.